OBSERVATIONS OF FORMIC ACID TOWARD GALACTIC HOT MOLECULAR CORES

\underline{SHENG-YUAN LIU}, DAVID M. MEHRINGER, and LEWIS E. SNYDER, Department of Astronomy, University of Illinois, 1002 W.Green, Urbana, IL 61801.

Formic acid shares structural elements with the common interstellar molecule methyl formate (HCOOCH_{3}) and the elusive biomolecule acetic acid (CH_{3}COOH). It is the simplest organic acid, but it has only been identified in astronomical sources by fairly weak lines in single-element telescope surveys. In order to study the successive formation of biomolecules in the ISM, we carried out a survey of HCOOH toward galactic hot molecular cores with the Berkeley-Illinois-Maryland Association (BIMA) Array. Up to five transitions with rest frequencies near 87 GHz have been detected simultaneously in three sources: Sgr B2(N), Orion, and W51. HCOOH was found to have an excitation temperature of at least 100 K. The formation of HCOOH is probably related to grain-surface chemistry in hot cores. Abundance comparisons with other molecules such as HCOOCH_{3} and ethyl cyanide (C_{2}H_{5}CN) will be discussed.

This work was partially funded by NSF AST96 -13999 and the University of Illinois.